OpenStudy (anonymous):

simplify sq root of 40 and 200

OpenStudy (anonymous):

Square root of 40 is 2 root 10 and square root of 200 is 10 root 2

OpenStudy (anonymous):

does "and" mean add, multiply or are they two different problems?

OpenStudy (anonymous):

6.324 and 14.14

OpenStudy (anonymous):

i assume when it says "simplify" it means to write in simplest radical form. in which case lagrange has it

OpenStudy (anonymous):

\[\sqrt{20}=\sqrt{4\times 10}=\sqrt{4}\times \sqrt{10}=2\sqrt{10}\]

OpenStudy (anonymous):

it probably does not mean give a decimal approximation

OpenStudy (anonymous):

my mistake..sorry

OpenStudy (anonymous):

likewise \[\sqrt{200}=\sqrt{100\times 2}=\sqrt{100}\times \sqrt{2}=10\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{40}= \sqrt{2 x 2 x 2 x 5} = 2 \sqrt{10}\] \[\sqrt{200}= \sqrt{2x2x2x5x5}\]